OpenStudy (cinnamonbun):

What would the surface area of a cube be if each dimension were multiplied by the scale factor 1/3? One side is 12 meters

OpenStudy (cinnamonbun):



OpenStudy (welshfella):

the area would be (1/3)^2 of the original

OpenStudy (cinnamonbun):

@welshfella 1/9?

OpenStudy (erak):

Yes, for example the SA for the 12m cube is SA = 6s^2 SA = 6(12)^2 SA = 864 And if you took the side and applied the scale factor, 12*(1/3) = 4m SA = 6s^2 SA = 6(4)^2 SA = 96 And if you found the ratio: 96:864 = 1:9 or 1/9